OpenStudy (anonymous):

Solve the system of equations and give the value x: 4x + y + -14 3x + 4y = -30

OpenStudy (anonymous):

12x + 3y = -42 12x + 16y = -120 ----------------- 12x + 3y = -42 -12x - 16y = 120 ----------------- -13y = 78 y = -6 x = 2

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

Yeah thanks

OpenStudy (anonymous):

yw